### v2

# VC-Density in Pairs of Ordered Vector Spaces

Abstract.
We show that the VC-density of any partitioned formula in a pair of ordered vector spaces is bounded above by twice the number of parameter variables. We also show that this bound is optimal and, as a by-product, we prove that no dense pair of o-minimal structures is dp-minimal.
1. Introduction
We study pairs of ordered vector spaces over an ordered field, focusing on the VC-densities of formulas. The concept of VC-dimension is defined in the context of statistical learning by Vapnik and Chervonenkis; hence the acronym VC. A finer invariant is the VC-density of a formula. We define these concepts in the last section, along with some facts about them.
Let $V$ be an ordered vector space over an ordered field $K$, and let $W$ be a proper dense subspace of $V$. We consider $V$ and $W$ as structures in the language $L_{\operatorname{ovs}}:=\{+,-,0,1,<\}\cup\{\ell_{\lambda}:\lambda\in K\}$. Then $W$ is an elementary substructure of $V$. Also, let $L_{\operatorname{ovs}}^{d}:=L_{\operatorname{ovs}}\cup\{U\}$ be the enrichment of $L_{\operatorname{ovs}}$ by a unary relation symbol $U$. Letting $(V,W)$ denote the $L_{\operatorname{ovs}}^{d}$-structure where $U$ is interpreted as $W$, our main result is as follows.
Theorem 1.1.
The VC-density in $(V,W)$ of a partitioned $L_{\operatorname{ovs}}^{d}$-formula $\phi(x;y)$ is at most $2|y|$.
We prove that this is the best possible bound; more precisely, we show that there is a partitioned $L_{\operatorname{ovs}}^{d}$-formula $\phi(x;y)$, where both $x$ and $y$ are single variables such that the VC-density of $\phi$ is $2$. Actually, that formula allows us to show that no dense pair of o-minimal structures is dp-minimal.
The paper [1] relates VC-density with uniform definability of types over finite sets as defined by Guingona in [6]; we restate this relation in the last section as Theorem 4.4. In the light of that, we actually prove the following stronger result.
Theorem 1.2.
The theory of dense pairs of ordered vector spaces over an ordered field $K$ has the VC 2 property.
The VC 2 property is a stronger version of the uniform definability of types over finite sets; we introduce it in the last section of the paper.
The structure $(V,W)$ is an example of dense pair of o-minimal structures as defined in [3]. We recall the definition of this concept in Section 3, and we also mention some facts about dense pairs of o-minimal structures. One of those facts is a quantifier elimination result as in the first part of Theorem 3.1 below. We prove a quite different description of definable families of subsets of $V$. In order to explain that, we introduce two kinds of definable sets. For two functions $a,b:V^{l}\to V$ definable in the pair $(V,W)$ such that $a(t)<b(t)$ for all $t\in V^{l}$, we define
$Q_{a,b}:=\left\{(x,y)\in V\times V^{l}:x\in[a(y),b(y)]\right\},$
and for $g:V^{l}\to V$ definable in the vector space $V$, we let
$C_{g}:=\left\{(x,y)\in V\times V^{l}:x\in g(y)+W\right\}.$
Theorem 1.3.
Let $Z\subseteq V^{1+l}$ be definable in the pair $(V,W)$. Then $Z$ is a boolean combination of sets varying in the following list:
  1. (1)
    Subsets of $V^{1+l}$ that are definable in the vector space $V$.
  2. (2)
    The sets $Q_{a,b}$ for $a,b$ as above.
  3. (3)
    Sets of the form $V\times Y$, where $Y\subseteq V^{l}$ is definable in the pair $(V,W)$.
  4. (4)
    The sets $C_{g}$ for $g$ as above.
Notations and Conventions. The set $\mathbb{N}$ of natural numbers contains 00 and +=$\mathbb{N}_{+}=\mathbb{N}\setminus\{0\}$. We let the letters $m,n,k,l$ vary in $\mathbb{Z}$, and if $m$ is in $\mathbb{N}$ (or $\mathbb{N}_{+}$), we simply write $m\geq 0$ (or $m>0$).
We let the letters $x,y,z$ denote tuples of variables, possibly with some decorations. In the same way, we do not use separate notations for elements and tuples of elements of sets. We make sure that this does not cause any confusion. One way to do this is to express the lengths of tuples using absolute value notation.
For a first-order language $L$, an $L$-structure $\mathcal{M}$, and $A\subseteq M$, we say that a subset of $M^{n}$ is $L(A)$-definable if it is defined by a formula of the form $\phi(x,a)$, where $\phi$ is an $L$-formula and $a$ is a tuple of elements of $A$; we refer to such a formula as an $L(A)$-formula. We write $L$-definable and $L$-formula instead of $L(\emptyset)$-definable and $L(\emptyset)$-formula. We also write definable in $\mathcal{M}$ in the place of $L(M)$-definable.
2. Semilinear Geometry
Let $K$ be an ordered field, and let $L_{\operatorname{ovs}}$ be the language defined in the Introduction. Also let $T_{\operatorname{ovs}}$ be the $L_{\operatorname{ovs}}$-theory of ordered vector spaces over $K$ with a distinguished positive element $1$. It is well-known that $T_{\operatorname{ovs}}$ is an o-minimal theory.
Letting $x=(x_{1},\dots,x_{n})$ and $\lambda=(\lambda_{1},\dots,\lambda_{n})\in K^{n}$, we use the dot-product notation $\lambda x$ to denote $\sum_{i=1}^{n}\ell_{\lambda_{i}}(x_{i})$.
Fix a model $V$ of $T_{\operatorname{ovs}}$. Before proving some results on the sets definable in the vector space $V,$ we recall some well-known results on them; we refer the reader to the last section of Chapter 1 in the book [4] by van den Dries for the details. (There, the language does not contain $1$, but that is immaterial for our discussions here.) We use the dot-product notation for tuples of elements of $V$ as well.
A function $f:V^{n}\to V$ is called affine if $f(x)=a+\lambda x$ for some $a\in V$ and $\lambda\in K^{n}$; if $a\in A$ for some subset $A$ of $V$, then we say that $f$ is defined over $A$. A basic semilinear set is a set of the form
(2.1) $\big{\{}v\in V^{n}:f_{1}(v)=\dots=f_{r}(v)=0,g_{1}(v)<0,\dots,g_{s}(v)<0\big{\}},$
where $f_{1},\dots,f_{r},g_{1},\dots,g_{s}$ are affine. If $f_{1},\dots,f_{r},g_{1},\dots,g_{s}$ are defined over $A$, then we say that this set is basic $A$-semilinear. An $A$-semilinear set is the union of finitely many basic $A$-semilinear sets.
The $L_{\operatorname{ovs}}(A)$-definable sets in $V$ are precisely the $A$-semilinear sets. So we use these notions interchangeably. Similarly, we use $A$-semilinear function in the place of ‘$L_{\operatorname{ovs}}(A)$-definable function in $V$’. When we do not wish to specify the parameter set, we write (basic) semilinear in the place of (basic) $V$-semilinear.
It is well-known that $A$-semilinear functions are locally affine, in the sense that there is a decomposition of the domain into finitely many $A$-semilinear sets such that the function agrees with an affine function defined over $A$ on each of those semilinear sets.
If $X\subseteq V^{n+l}$ and $t\in V^{l}$, then we let $X(t)$ denote the section of $X$ over $t$:
$x\in X(t)\Longleftrightarrow(x,t)\in X.$
For the rest of this section, we also fix a dense subspace $W$ of $V$. Considered as an $L_{\operatorname{ovs}}$-structure, $W$ is an elementary substructure of $V$.
Lemma 2.1.
Let $S\subseteq V^{n}$ be a basic semilinear set such that $S\cap W^{n}\neq\emptyset$, and let $s\in S$. Then there is a basic $W$-semilinear subset $T$ of $S$ containing $s$.
Proof.
We proceed by induction on $n$.
If $n=1$, then either $S=\{s\}$ or $S=(a,b)$ for some $a,b\in V\cup\{\pm\infty\}$. In the first case, $s\in W$ and hence $S$ is basic $W$-semilinear. In the second case, take $w_{1},w_{2}\in W$ such that $a<w_{1}<s<w_{2}<b$, and let $T=(w_{1},w_{2})$.
Now assume that the result holds for $n$ and let’s prove it for $n+1$. So, let $S\subseteq V^{n+1}$ be a basic semilinear set that intersects $W^{n+1}$ and let $s\in S$. If $S$ is open, then we could take $T$ to be an open box in $S$ containing $s$ whose vertices are in $W$. If $S$ is not open, then, after changing the order of coordinates, we have that $S$ is the graph of the restriction of an affine $f:V^{n}\to V$ to a basic semilinear $S^{\prime}\subseteq V^{n}$; write $s=(s^{\prime},f(s^{\prime}))$, where $s^{\prime}\in S^{\prime}$. Since $S\cap W^{n+1}\neq\emptyset$, the function $f$ is defined over $W$. Using the induction hypothesis, take a basic $W$-semilinear $T^{\prime}\subseteq S^{\prime}$ containing $s^{\prime}$. Now $s$ is contained in the graph of the restriction of $f$ to $T^{\prime}$. We are done by taking $T$ as that graph. ∎
Proposition 2.2.
Let $S\subseteq V^{n}$ be a basic semilinear set such that $S\cap W^{n}\neq\emptyset$ and let $f:V^{n}\rightarrow V$ be an affine function. Then
$f(S\cap W^{n})=(f(0)+W)\cap f(S).$
Proof.
It is clear that $f(S\cap W^{n})\subseteq(f(0)+W)\cap f(S)$. For the other inclusion, let $s\in S$ such that $f(s)-f(0)\in W$. Our aim is to show that $N\cap S\cap W^{n}\neq\emptyset$, where $N:=f^{-1}(f(s))$. Even though $f$ is not necessarily defined over $W$, $N$ is a basic $W$-semilinear set. Using Lemma 2.1, take a basic $W$-semilinear subset $T$ of $S$ containing $s$. Then $N\cap T$ is an $L_{\operatorname{ovs}}(W)$-definable set and $s\in N\cap T$. Therefore, we have $N\cap T\cap W^{n}\neq\emptyset$ by Tarski-Vaught test. Hence $N\cap S\cap W^{n}\neq\emptyset$. ∎
Since semilinear maps are piecewise affine, we get the following consequence of the proposition above.
Proposition 2.3.
Let $S\subseteq V^{n+l}$ and $f:V^{n+l}\to V$ be $A$-semilinear. Then there are $A$-semilinear sets $T_{1},\dots,T_{r}\subseteq V^{1+l}$ and affine $g_{1},\dots,g_{r}:V^{l}\to V$ defined over $A$ such that for every $t\in V^{l}$ we have
$f(S(t)\cap W^{n},t)=\bigcup_{i=1}^{r}(g_{i}(t)+W)\cap T_{i}(t).$
Proof.
Let $U_{1},\dots,U_{r}\subseteq V^{n+l}$ be $A$-semilinear and let $h_{1},\dots,h_{r}:V^{n+l}\to V$ be affine functions defined over $A$ such that $V^{n+l}=U_{1}\cup\dots\cup U_{r}$ and $f|_{U_{i}}=h_{i}|_{U_{i}}$ for all $i\in\{1,\dots,r\}$.
For $i\in\{1,\dots,r\}$ put $S_{i}:=U_{i}\cap S$. Then for every $t\in V^{l}$ we have
$f(S(t)\cap W^{n},t)$ $=f(\bigcup_{i=1}^{r}S_{i}(t)\cap W^{n},t)$
$=\bigcup_{i=1}^{r}f(S_{i}(t)\cap W^{n},t)$
$=\bigcup_{i=1}^{r}h_{i}(S_{i}(t)\cap W^{n},t)$
We may assume that $S_{i}$ are basic $A$-semilinear after writing them as a union of basic $A$-semilinear sets. Then using Proposition 2.2, for all $i$ and all $t\in V^{l}$, we get
$h_{i}(S_{i}(t)\cap W^{n},t)=(h_{i}(0,t)+W)\cap h_{i}(S_{i}(t),t).$
Clearly, for each $i$, there is an $A$-semilinear set $T_{i}\subseteq V^{1+l}$ such that $T_{i}(t)=h_{i}(S_{i}(t),t)$ for all $t\in V^{l}$. Letting $g_{i}:V^{l}\to V$ be defined as $g_{i}(x)=h_{i}(0,x)$, we obtain the desired conclusion. ∎
3. Definable Families in Dense Pairs of Ordered Vector Spaces
Let $L\supseteq\{+,-,0,1,<\}$ be a language and $T$ a complete o-minimal $L$-theory extending the theory of ordered abelian groups with a distinguished positive element $1$. Let $L^{d}:=L\cup\{U\}$, where $U$ is a new unary relation symbol. A dense pair of models of $T$ is an $L^{d}$-structure $(\mathcal{M},\dots,N)$, where $\mathcal{M}\models T$, $N$ is the universe of an $L^{d}$-structure $\mathcal{N}$, which is a proper elementary substructure of $\mathcal{M}$, and $N$ is dense in $M$. We denote such a dense pair simply as $(\mathcal{M},\mathcal{N})$. When $L$ is $L_{\operatorname{ovs}}$, we keep the terminology from the previous section and use ‘semilinear’ interchangeably with ‘$L(M)$-definable’.
It is clear that being a dense pair is expressible in $L^{d}$ and it is shown in [3] that the $L^{d}$-theory of such pairs is indeed complete; we denote that theory by $T^{d}$. The following results on definable sets are by-products of the proof of completeness.
Theorem 3.1.
Let $(\mathcal{M},\mathcal{N})$ be a model of $T^{d}$.
  1. (1)
    Every $L^{d}$-definable subset of $M^{n}$ can be defined by a boolean combination of formulas of the form
    $\exists y_{1}\cdots\exists y_{m}(U(y_{1})\wedge\dots\wedge U(y_{m})\wedge\phi(x_{1},\dots,x_{n},y_{1},\dots,y_{m})),$
    where $\phi$ is an $L$-formula.
  2. (2)
    A set $X\subseteq N^{n}$ is $L^{d}(M)$-definable if and only if there is an $L(M)$-definable subset $S$ of $M^{n}$ such that $X=S\cap N^{n}$.
Below we prove a strengthening of the second part of this theorem for models of $T_{\operatorname{ovs}}^{d}$; namely a version that is uniform on definable families.
The following is a direct consequence of Corollary 2.7 of [3].
Lemma 3.2.
Let $\mathcal{M},\mathcal{M}^{\prime},\mathcal{N}$ be models of $T$ such that $\mathcal{N}\subseteq\mathcal{M}^{\prime}\subseteq\mathcal{M}$ and both $(\mathcal{M},\mathcal{N})$ and $(\mathcal{M}^{{}^{\prime}},\mathcal{N})$ are models of $T^{d}$. Then $(\mathcal{M}^{{}^{\prime}},\mathcal{N})\preceq(\mathcal{M},\mathcal{N})$.
The next result is proved in [5] in a more general setting than dense pairs of o-minimal structures.
Theorem 3.3 (Lemma 3.3 of [5]).
Let $(\mathcal{M},\mathcal{N})$ be a model of $T^{d}$, and let $Z\subseteq M^{1+l}$ be $L^{d}(A)$-definable for some $A\subseteq M$. Then there are $m,n,p>0$ and for each $i=1,\ldots,m$ there are
  • $L^{d}(A)$-definable subsets $X_{i1},\dots,X_{ip}$ of $M^{n+l}$ such that $X_{ij}(t)\subseteq N^{n}$ for all $t\in M^{l}$,
  • $L(A)$-definable functions $h_{i1},\ldots,h_{ip}:M^{n+l}\rightarrow M$,
  • an $L^{d}(A)$-definable function $a_{i}:M^{l}\rightarrow M\cup\{\infty\}$,
such that for $t\in M^{l}$,
  1. i.
    $-\infty=a_{0}(t)\leq a_{1}(t)\leq\dots\leq a_{m}(t)=\infty$ is a decomposition of $M$, and
  2. ii.
    one of the following holds:
    1. (a)
      $[a_{i-1}(t),a_{i}(t)]\cap Z(t)=P_{i}(t)$ or
    2. (b)
      $[a_{i-1}(t),a_{i}(t)]\cap Z(t)=[a_{i-1}(t),a_{i}(t)]\setminus P_{i}(t)$,
where $P_{i}(t)=\bigcup_{j=1}^{p}h_{ij}(X_{ij}(t),t)$.
Now we focus on the case of ordered vector spaces; so $T$ is $T_{\operatorname{ovs}}$. We denote models of $T_{\operatorname{ovs}}^{d}$ as $(V,W)$. We first prove a version of the second part of Theorem 3.1 that takes the parameters into account.
Lemma 3.4.
Let $(V,W)$ be a model of $T_{\operatorname{ovs}}^{d}$, and let $X\subseteq W^{n}$ be $L^{d}_{\operatorname{ovs}}(A)$-definable. Then $X=S\cap W^{n}$ for an $A$-semilinear set $S\subseteq V^{n}$.
Proof.
Let $\operatorname{dcl}$ denote the definable closure in $V$ and let $V^{{}^{\prime}}=\operatorname{dcl}(A\cup W)$. Clearly, $V^{\prime}$ is a model of $T_{\operatorname{ovs}}$ and hence $V^{\prime}\preceq V$.
If $V^{\prime}=W$, then $A\subseteq W$. Then the conclusion easily follows from the first part of Theorem 3.1.
So we may assume $W\neq V^{\prime}$ and thus $(V^{\prime},W)$ is a model of $T_{\operatorname{ovs}}^{d}$. Moreover by Lemma 3.2, we have $(V^{\prime},W)\preceq(V,W)$. As $A\subseteq V^{\prime}$, the set $X(V^{\prime}):=X\cap(V^{\prime})^{n}$ is definable in $(V^{\prime},W)$ and $X(V^{\prime})\subseteq W^{n}$.
Now applying part (2) of Theorem 3.1, there is a semilinear subset $S^{{}^{\prime}}$ of $(V^{{}^{\prime}})^{n}$ such that
$S^{{}^{\prime}}\cap W^{n}=X(V^{{}^{\prime}}).$
It is clear that $S^{\prime}$ is indeed $L_{\operatorname{ovs}}(A)$-definable. So let $\phi(x)$ be an $L_{\operatorname{ovs}}(A)$-formula defining $S^{\prime}$ and let $S$ be the set defined in $V$ by $\phi(x)$. Since $(V^{\prime},W)\preceq(V,W)$, we get that
$S\cap W^{n}=X.$
The result below is the promised strengthening of part (2) of Theorem 3.1.
Proposition 3.5.
Let $(V,W)$ be a model of $T_{\operatorname{ovs}}^{d}$, and let $X\subseteq V^{n+l}$ be $L_{\operatorname{ovs}}^{d}(A)$-definable such that $X(t)\subseteq W^{n}$ for any $t\in V^{l}$. Then there are $A$-semilinear sets $S_{1},\ldots,S_{k}\subseteq V^{n+l}$ such that for each $t\in V^{l}$, there is $i\in\{1,\dots,k\}$ with $X(t)=S_{i}(t)\cap W^{n}$.
Proof.
Given an $L_{\operatorname{ovs}}(A)$-definable $S\subseteq V^{n+l}$, let $\phi_{S}(y)$ denote the formula expressing $X(y)\neq S(y)\cap W^{n}$ and consider the following set of formulas:
[…]
If the conclusion of the proposition does not hold, then this set is finitely satisfiable in $(V,W)$. Therefore there is an elementary extension $(V^{*},W^{*})$ of $(V,W)$ and $t^{*}\in(V^{*})^{l}$ such that
$X^{*}(t^{*})\neq S^{*}(t^{*})\cap(W^{*})^{n},$
for any $L_{\operatorname{ovs}}(A)$-definable $S$, where $X^{*}$ and $S^{*}$ denote the realizations of $X$ and $S$ in $(V^{*},W^{*})$. However, applying Lemma 3.4 to $X^{*}(t^{*})$, there is an $L_{\operatorname{ovs}}(A\cup\{t^{*}\})$-definable set $T^{*}\subseteq(V^{*})^{n}$ such that $X^{*}(t^{*})=T^{*}\cap(W^{*})^{n}$ and such $T^{*}$ is of the form $S^{*}(t^{*})$ for some $L_{\operatorname{ovs}}(A)$-definable $S\subseteq V^{n+l}$. This contradiction finishes the proof. ∎
Proposition 3.6.
Let $(V,W)$ be a model of $T_{\operatorname{ovs}}^{d}$. Let $h:V^{n+l}\rightarrow V$ be $A$-semilinear and let $X\subseteq V^{n+l}$ be $L_{\operatorname{ovs}}^{d}(A)$-definable such that $X(t)\subseteq W^{n}$ for each $t\in V^{l}$. Then there are $L_{\operatorname{ovs}}^{d}(A)$-definable $Y_{1},\dots,Y_{k}\subseteq V^{l}$, and for each $i\in\{1,\dots,k\}$, there are affine functions $g_{i1},\ldots,g_{iq_{i}}:V^{l}\rightarrow V$ defined over $A$, and $A$-semilinear sets $T_{i1},\ldots,T_{iq_{i}}\subseteq V^{1+l}$ such that for each $t\in Y_{i}$ we have
$h(X(t),t)=\bigcup_{j=1}^{q_{i}}(g_{ij}(t)+W)\cap T_{ij}(t).$
Proof.
Let $S_{1},\dots,S_{k}\subseteq V^{n+l}$ be as in Proposition 3.5. Define $Y_{0}=\emptyset$ and $Y_{i}:=\{t\in V^{l}:X(t)=S_{i}(t)\cap W^{n}\}\setminus(Y_{0}\cup\dots\cup Y_{i-1})$ for $i>0$. Then $Y_{1},\dots,Y_{k}$ are $L_{\operatorname{ovs}}^{d}(A)$-definable, and the nonempty members of this collection form a partition of $V^{l}$.
Now using Proposition 2.3, we may choose, for each $i$, the affine functions $g_{i1},\ldots,g_{iq_{i}}$ and semilinear sets $T_{i1},\ldots,T_{iq_{i}}$ satisfying the requirements.
We put this result together with Theorem 3.3 to prove Theorem 1.3 from the Introduction; this version is slightly stronger as we take parameters into account. The sets $Q_{a,b}$ and $C_{g}$ are as defined in the Introduction.
Theorem 3.7.
Let $Z\subseteq V^{1+l}$ be $L_{\operatorname{ovs}}^{d}(A)$-definable in a model $(V,W)$ of $T_{\operatorname{ovs}}^{d}$. Then $Z$ is a boolean combination of sets varying in the following list:
  1. (1)
    $A$-semilinear subsets of $V^{1+l}$.
  2. (2)
    The sets $Q_{a,b}$ for $L_{\operatorname{ovs}}^{d}(A)$-definable functions $a,b$ such that $a(t)<b(t)$ for all $t\in V^{l}$.
  3. (3)
    Sets of the form $V\times Y$, where $Y\subseteq V^{l}$ is $L_{\operatorname{ovs}}^{d}(A)$-definable.
  4. (4)
    The sets $C_{g}$ for affine functions $g$ defined over $A$.
Proof.
For $i\in\{1,\dots,m\},j\in\{1,\dots,p\}$, let $X_{ij}\subseteq V^{n+l}$, $h_{ij}:V^{n+l}\rightarrow V$, $a_{i}:V^{l}\rightarrow V\cup\{\infty\}$, and $P_{i}\subseteq V^{1+l}$ be as in Theorem 3.3. Put $Z_{i}:=Z\cap Q_{a_{i-1},a_{i}}$ and $R_{i}:=\{y\in V^{l}:Z_{i}(y)=P_{i}(y)\}$, a definable subset of $V^{l}$. Then
$Z=\bigcup_{i=1}^{m}Z_{i}=\bigcup_{i=1}^{m}Z_{i}^{+}\cup Z_{i}^{-},$
where $Z_{i}^{+}=Z_{i}\cap(V\times R_{i})$ and $Z_{i}^{-}=Z_{i}\cap(V\times(V^{l}\setminus R_{i}))$. So for our purposes, it is enough to show that both $Z_{i}^{+}$ and $Z_{i}^{-}$ are boolean combinations of sets in the list.
Write
$Z_{i}^{+}$ $=\bigcup_{y\in R_{i}}Z_{i}(y)\times\{y\}$
$=\bigcup_{y\in R_{i}}P_{i}(y)\times\{y\}$
$=\bigcup_{y\in R_{i}}(\bigcup_{j=1}^{p}h_{ij}(X_{ij}(y),y))\times\{y\}$
$=\bigcup_{j=1}^{p}(\bigcup_{y\in R_{i}}h_{ij}(X_{ij}(y),y)\times\{y\})$
$=\bigcup_{j=1}^{p}\big{(}(\bigcup_{y\in V^{l}}h_{ij}(X_{ij}(y),y)\times\{y\})\cap(V\times R_{i})\big{)}.$
Therefore it suffices to show that $Z_{ij}^{+}:=\bigcup_{y\in V^{l}}h_{ij}(X_{ij}(y),y)\times\{y\}$ is of the required form for each $j$. Using Proposition 3.6, take $L_{\operatorname{ovs}}^{d}(A)$-definable $Y_{ijk}\subseteq V^{l}$ for $k=1,\dots,r$ and for each $k\in\{1,\dots,r\}$, affine functions $g_{ijk1},\ldots,g_{ijkq_{k}}:V^{l}\rightarrow V$ defined over $A$, $A$-semilinear sets $T_{ijk1},\ldots,T_{ijkq_{k}}\subseteq V^{1+l}$ such that for each $y\in Y_{ijk}$ we have
$h_{ij}(X_{ij}(y),y)=\bigcup_{l=1}^{q_{k}}(g_{ijkl}(y)+W)\cap T_{ijkl}(y).$
So
$Z_{ij}^{+}=\bigcup_{k=1}^{r}\bigcup_{y\in Y_{ijk}}(\bigcup_{l=1}^{q_{k}}(g_{ijkl}(y)+W)\cap T_{ijkl}(y))\times\{y\},$
and hence it is enough to show that for each $i,j,k,l$, the set
$\bigcup_{y\in Y_{ijk}}\big{(}(g_{ijkl}(y)+W)\cap T_{ijkl}(y)\big{)}\times\{y\}$
is of the required form. It indeed is, since it equals
$C_{g_{ijkl}}\cap T_{ijkl}\cap(V\times Y_{ijk}).$
As for $Z_{i}^{-}$, one may show that that
$Z_{i}^{-}=(Q_{a_{i-1},a_{i}}\setminus P_{i})\cap(V\times(V\setminus R_{i})).$
Hence going through similar steps, we see that $Z_{i}^{-}$ is also of the required form. ∎
4. VC-density and UDTFS
Below $\mathcal{M}$ is a structure in a language $L$ and its theory is denoted by $T$. We deal with partitioned $L$-formulas $\phi(x;y)$, which means that the set of free variables of $\phi$ are partitioned into two: object variables $x=(x_{1},\dots,x_{m})$ and parameter variables $y=(y_{1},\dots,y_{n})$.
For a subset $A\subseteq M^{m}$, we let $A\cap\phi(\mathcal{M};y)$ denote the collection
$\big{\{}A\cap\phi(\mathcal{M};b):b\in M^{n}\big{\}}.$
If, for instance, $A$ is finite with $k$ elements, then $A\cap\phi(\mathcal{M};y)$ has at most $2^{k}$ members.
The shatter function $\pi_{\phi}:\mathbb{N}\to\mathbb{N}$ of $\phi(x;y)$ is defined as follows:
$\pi_{\phi}(k):=\max\big{\{}|A\cap\phi(\mathcal{M};y)|:A\subseteq M^{m},|A|=k\big{\}}.$
If there is $l\geq 0$ with $\pi_{\phi}(l)<2^{l}$, then the VC-dimension of $\phi(x;y)$ is the largest $k$ such that $\pi_{\phi}(k)=2^{k}$. Otherwise we say that the VC-dimension of $\phi(x;y)$ is infinity.
A result proved independently by Sauer (in [8]) and Shelah (in [9]) states that if the VC-dimension of $\phi(x;y)$ is finite, say $d$, then the shatter function $\pi_{\phi}$ is bounded by a polynomial of $k$ of degree $d$. This allows us to define the VC-density of $\phi(x;y)$ to be the infimum of the real numbers $r\in\mathbb{R}$ such that $\pi_{\phi}(k)=O(k^{r})$.
We do not actually calculate the VC-density of formulas in the dense pairs of ordered vector spaces. We use its relation with the concept of the definability of types as defined below.
Let $\Delta(x;y)$ be a finite set of partitioned $L$-formulas; let $m=|x|$ and $n=|y|$. Given a subset $B\subseteq M^{n}$, a set of $L$-formulas that are either of the form $\phi(x;b)$ or of the form $\neg\phi(x;b)$, where $\phi\in\Delta$ and $b\in B$ is called a $\Delta$-type over $B$ if it is maximally consistent in $T$. (In many sources this is called a complete $\Delta$-type.) We denote the set of $\Delta$-types over $B$ as $S^{\Delta}(B)$.
Definition 4.1.
A finite set $\Delta(x;y)$ has uniform definability of types over finite sets (UDTFS) with $d$ parameters (in $\mathcal{M}$) if there are collections
$(\phi^{1}(y;y_{1},\dots,y_{d}))_{\phi\in\Delta},\dots,(\phi^{r}(y;y_{1},\dots,y_{d}))_{\phi\in\Delta}$
of $L$-formulas such that for every finite $B\subseteq M^{n}$ and $p\in S^{\Delta}(B)$ there are $i\in\{1,\dots,r\}$ and $b_{1},\dots,b_{d}\in B$ such that for every $b\in B$
$\phi(x;b)\in p\Longleftrightarrow\mathcal{M}\models\phi^{i}(b;b_{1},\dots,b_{d}).$
The collections $(\phi^{1}(y;y_{1},\dots,y_{d}))_{\phi\in\Delta},\dots,(\phi^{r}(y;y_{1},\dots,y_{d}))_{\phi\in\Delta}$ as in this definition are said to define $\Delta$.
It is shown in [2] that in any structure with NIP, every singleton $\{\phi(x;y)\}$ has UDTFS with $d$ parameters for some $d$. Using Corollary 3.2 of [7], the same holds for any dense pair of models of an o-minimal theory extending the theory of divisible ordered abelian groups as in Section 3. We are aiming to show that in a pair $(V,W)$ of models of $T_{\operatorname{ovs}}$, any finite $\Delta(x;y)$ with a single object variable actually has UDTFS with $2$ parameters. While doing this, we use a few results from [1]. We start with the following, which is Corollary 5.6 in [1].
Theorem 4.2.
Let $\Phi$ be a set of partitioned formulas in a single object variable $x$ such that
  • every partitioned formula with object variable $x$ is equivalent in $T$ to a boolean combination of formulas from $\Phi$.
  • every finite subset of $\Phi$ has UDTFS with $d$ parameters.
Then every finite set of $L$-formulas with a single object variable has UDTFS with $d$ parameters.
In [1], a structure satisfying the conclusion of this theorem is said to have the VC d property; we will be using this terminology as well. Another result from [1] is that every weakly o-minimal structure has the VC 1 property; Theorem 6.1.
The following is a consequence of Lemma 5.2 of [1]. We present a proof of it, so that we do not have to introduce even more notation from [1].
Proposition 4.3.
Let $\mathcal{M}=(M,<,\dots)$ be an expansion of a dense linear order without endpoints. Let $\Phi(x,y)$ be the set of formulas of the form $x<g(y)$, where $x$ is a single variable, $y$ is a variable of various length, and $g$ is an $\emptyset$-definable function. Then every finite subset of $\Phi$ has UDTFS with a single parameter in $\mathcal{M}$.
Proof.
Let $\Delta$ be a finite subset of $\Phi$; so we may assume that the length of $y$ is fixed, say $n$. Let $g_{1}(y),\dots,g_{k}(y)$ be $\emptyset$-definable functions such that elements of $\Delta$ are $x<g_{i}(y)$ for $i=1,\dots,k$.
We claim that the collections $(g_{j}(y_{1})\leq g_{i}(y))_{i=1,\dots,k}$ for $j=1,\dots,k$ and $(y\neq y)_{i=1,\dots,k}$ define every $\Delta$-type over any finite set $B$.
Let $B$ be a finite subset of $M^{n}$ and $q\in S^{\Delta}(B)$. If $q$ does not contain any of the formulas $x<g_{i}(b)$, then it is defined by $y\neq y$. So assume that $q$ contains a formula of the form $x<g_{i}(b)$ and let $g_{i_{1}}(b_{1})$ be the smallest element of the set $\{g_{i}(b):x<g_{i}(b)\text{ is in }q\}$. Then for any $b\in B$, we have $x<g_{i}(b)$ is in $q$ if and only if $\mathcal{M}\models g_{i_{1}}(b_{1})\leq g_{i}(b)$. Therefore $(g_{i_{1}}(b_{1})\leq g_{i}(y))_{i=1,\dots,k}$ defines $q$. ∎
Next we state the relation of the definability of types and VC-density as given in [1].
Theorem 4.4 (Corollary 5.8 of [1]).
Suppose that $\mathcal{M}$ has the VC d property. Then the VC-density of every formula $\phi(x;y)$ is bounded by $d|y|$.
In this theorem, $x$ is of arbitrary length.
Now we are in a position to prove Theorem 1.2, and Theorem 1.1 is a consequence of it combined with Theorem 4.4 above.
Theorem 4.5.
Any model $(V,W)$ of $T_{\operatorname{ovs}}^{d}$ has the VC 2 property.
Proof.
Fix a single variable $x$, and let $y$ vary among tuples of variables of various lengths.
Let $\Phi_{1}$ be the collection of partitioned $L_{\operatorname{ovs}}^{d}$-formulas consisting of all the formulas of the form $x<a(y)$, where $a$ is an $L_{\operatorname{ovs}}^{d}$-definable function. Let $\Phi_{2}$ be the set of $L_{\operatorname{ovs}}^{d}$-formulas of the form $U\big{(}x-g(y)\big{)}$, where $g$ is an affine function defined over $\emptyset$. Finally, let $\Phi_{3}$ be the set of $L_{\operatorname{ovs}}^{d}$-formulas in variables $y$, considered as a partitioned formula in variables $(x;y)$.
Put $\Phi:=\Phi_{1}\cup\Phi_{2}\cup\Phi_{3}$. By Theorem 3.7 and the Cell Decomposition Theorem for o-minimal theories, every partitioned $L_{\operatorname{ovs}}^{d}$-formula with object variable $x$ is equivalent in $(V,W)$ to a boolean combination of elements of $\Phi$. Therefore by Theorem 4.2, it is enough to show that every finite subset of $\Phi$ has UDTFS with $2$ parameters. So let $\Delta$ be a finite subset of $\Phi$. Possibly after adding some dummy variables into some formulas in $\Delta$, we may assume that the parameter variables of elements of $\Delta$ is a fixed tuple $y$ of length $l$. For $i=1,2,3$, let $\Delta_{i}:=\Delta\cap\Phi_{i}$.
By Proposition 4.3, there are collections $(\phi^{(1)}(y;y_{1}))_{\phi\in\Delta_{1}},\ldots,(\phi^{(s)}(y;y_{1}))_{\phi\in\Delta_{1}}$ such that given finite $B\subseteq V^{l}$ and $q\in S^{\Delta_{1}}(B)$, there are $i\in\{1,\ldots,s\}$ and $b_{q}\in B$ such that for every $\phi\in\Delta_{1}$ and for every $b\in B$ we have
$\phi(x;b)\in q\Longleftrightarrow(V,W)\models\phi^{(i)}(b;b_{q}).$
Enumerate $\Delta_{2}$ as $\psi_{1}(x;y),\dots,\psi_{r}(x;y)$, where $\psi_{j}(x;y)$ is $U\big{(}x-g_{j}(y)\big{)}$ for some affine function $g_{j}:V^{l}\to V$ defined over $\emptyset$.
Let $k\in\{1,\dots,r+1\}$. For $i\in\{1,\dots,s\}$ and for $\phi\in\Delta_{1}$ define $\phi^{i,k}(y;y_{1},y_{2})$ as $\phi^{(i)}(y;y_{1})$. Now let $\psi_{j}$ be in $\Delta_{2}$. If $k\neq r+1$, define $\psi_{j}^{i,k}(y;y_{1},y_{2})$ to be
$U\big{(}g_{j}(y)-g_{k}(y_{2})\big{)}.$
We let $\psi_{j}^{i,r+1}(y;y_{1},y_{2})$ be $y\neq y$.
Note that for $\theta\in\Delta_{3}$ and given finite subset $B\subseteq V^{l}$, $q\in S^{\Delta}(B)$, and $b\in B$, we have
$\theta(x;b)\in q\Longleftrightarrow(V,W)\models\theta(b).$
So for $i\in\{1,\dots,s\}$ and $k\in\{1,\dots,r+1\}$, we define $\theta^{i,k}(y;y_{1},y_{2})$ as $\theta(y)$.
We claim that the collections $(\phi^{i,k})_{\phi\in\Delta}$ ($i=1,\dots,s$, $k=1,\dots,r+1$) define $\Delta$. So fix $B\subseteq V^{n}$ and $q\in S^{\Delta}(B)$. For $i=1,2,3$, let $q_{i}$ be the part of $q$ that arises from $\Delta_{i}$. Note that $q_{i}\in S^{\Delta_{i}}(B)$.
Take $b_{1}\in B$ and $i\in\{1,\ldots,s\}$ such that $\big{(}\phi^{(i)}(y;b_{1})\big{)}_{\phi\in\Delta_{1}}$ defines $q_{1}$. First suppose that there are $b_{2}\in B$ and $k\in\{1,\dots,r\}$ such that the formula $U(x-g_{k}(b_{2}))$ is in $q_{2}$. Then for all $b\in B$:
$\psi_{j}(x;b)\in q\Longleftrightarrow(V,W)\models\psi_{j}^{i,k}(b;b_{1},b_{2}).$
Therefore the collection $\big{(}\phi^{i,k}(y;b_{1},b_{2})\big{)}_{\phi\in\Delta}$ defines $q$ in this case.
Now suppose that $q_{2}$ does not contain a formula of the form $U(x-g_{j}(b))$. Then it is easy to see that $\big{(}\phi^{i,r+1}(y;b_{1},b_{2})\big{)}_{\phi\in\Delta}$ defines $q$ for any choice of $b_{2}\in B$.
Optimality of Theorem 1.1
We prove that our result is optimal by showing that the formula $y<x<y+1\vee U(x-y)$ has VC-density $2$. Note that this is a formula in the language of any dense pair of o-minimal structures; we actually prove the following general result.
Proposition 4.6.
Let $(\mathcal{M},\mathcal{N})$ be a dense pair of o-minimal structures, and let $\phi(x;y)$ be the formula $y\leq x<y+1\vee U(x-y)$. Then for any $k$, calculated in $(\mathcal{M},\mathcal{N})$, the quantity $\pi_{\phi}(k)$ is at least $\frac{k^{2}+k+2}{2}$.
Proof.
Let $A=\{a_{1},\dots,a_{k}\}$ where $a_{i}+1<a_{i+1}$ for each $i$ and $a_{i}-a_{j}\notin N$ for $i\neq j$. Clearly, $\phi(\mathcal{M},a_{i})\cap A=\{a_{i}\}$ for each $i$. Using the density of $N$, for each $i<j$, we may find $b$ such that $a_{i}\in[b,b+1)$ and $a_{j}\in b+N$. Therefore we have $\phi(\mathcal{M},b)\cap A=\{a_{i},a_{j}\}$. Hence the collection $A\cap\phi(\mathcal{M};y)$ contains all the singletons and two-element sets that are contained in $A$. Counting also the empty set, we get the desired lower bound. ∎
Putting this together with Theorem 1.1, we get the following.
Corollary 4.7.
Let $(V,W)$ be a dense pair of ordered vector spaces. Then the VC-density of $\phi$ is $2$.
For $n>0$, let $\phi_{n}(x;y_{1},\dots,y_{n})$ be the formula $\bigvee_{i=1}^{n}y_{i}\leq x<y_{i}+1\vee U(x-y_{i})$. One may modify the proof of the proposition above to show that $\pi_{\phi_{n}}(k)=O(k^{2n})$, hence we get that the VC-density of $\phi_{n}$ in any dense pair of ordered vector spaces is $2n$. Therefore, Theorem 1.1 is sharp for every possible tuple $y$ of variables.
Using the idea of the proof of the proposition above, we may show that no dense pair of o-minimal structures is dp-minimal. We use the definition of dp-minimality from [1].
Definition 4.8.
A theory $T$ is dp-minimal if for any model $\mathcal{M}$ of $T$ and formulas $\phi(x;y)$ and $\psi(x;y)$ with $|x|=1$, there are no sequences $(a_{i})_{i\in\mathbb{N}}$ and $(b_{i})_{i\in\mathbb{N}}$ from $M^{|y|}$ such that for any $i,j\in\mathbb{N}$ the following set of formulas has a realization in $\mathcal{M}$:
$\{\phi(x;a_{i}),\psi(x;b_{j})\}\cup\{\phi(x;a_{k}):k\neq i\}\cup\{\psi(x;b_{l}):l\neq j\}.$
Proposition 4.9.
Let $T$ be an o-minimal theory. Then $T^{d}$ is not dp-minimal.
Proof.
Let $(\mathcal{M},\mathcal{N})$ be a model of $T^{d}$. Let $\phi(x;y)$ and $\psi(x;y)$ be the formulas $y<x<y+1$ and $U(x-y)$. Take $(a_{i})_{i\in\mathbb{N}}$ from $M$ such that $a_{i}+1<a_{i+1}$ and $a_{i}-a_{j}\notin N$ for $i\neq j$. Also let $b_{i}=a_{i}$ for all $i\in\mathbb{N}$. Then given $i,j\in\mathbb{N}$, using the density of $N$ in $M$, we get $c\in(a_{i},a_{i}+1)\cap a_{j}+N$. It is clear that this $c$ is not in $\bigcup_{k\neq i,l\neq j}(a_{k},a_{k}+1)\cup a_{l}+N$. ∎